# Intertwined Orders and the Physics of High Temperature Superconductors

### v1

@@ II.1 Landau Fermi Liquids
The pole structure of the propagator implies that in the Fermi liquid there is a Fermi surface at $|\bm{p}|=k_{F}$ (here for simplicity we assumed an isotropic system) where the quasiparticles are long-lived. The analytic structure of the propagator also implies, among other things, that the resistivity has a $T^{2}$ temperature dependence. The existence of the Fermi surface is documented in different types of experimental probes such as quantum oscillations of resistivity, magnetization and electronic specific heat in magnetic fields, and in angle-resolved photoemission (ARPES) experiments.
@@ II.2 Ginzburg-Landau Description of the Superconducting State
The factor of $2$ in Eqs.(2), (3) and (4) anticipates the fact the superconducting order parameter is a condensate of Cooper pairs and, hence, has to transform as a charge $2$ complex field.
@@ II.3 BCS Theory
where $\hat{\Delta}(\bm{r})$ is the Fourier transform of the pair field operator of Eq.(5). This condition, known as off-diagonal long range order, implies that in the superconducting state the global $U(1)$ symmetry is spontaneously broken and that electric charge is conserved only module two.
@@ III High Temperature Superconductors are Different
Doping La2CuO4 with barium and/or strontium results in a relatively small number of mobile holes. The mobile holes drive the system into a superconducting state whose critical temperature $T_{c}$ has the shape of a dome as a function of doping with its maximum at about $15\%$ doping. Above the maximum superconducting $T_{c}$ is the strange metal phase and in-between there is the “pseudo-gap” regime. for large dopings, to the right of the superconducting dome there is a metallic phase widely believed to be a normal Fermi liquids. This pattern is present in the phase diagrams of all the copper-oxide superconductors, shown schematically in Fig.4.
@@
To this date there isn’t good theory of the strange metal phase and its explanation remains one of the challenges to Condensed Matter Theory. A commonly used theoretical proposal is to assume that there is a quantum critical point somewhere under the superconducting dome. Although this is an appealing proposal it has problems. One is that it is hard to find a quantum critical theory which naturally explains the linear resistivity. The other problem is that this behavior is seen not just in close proximity to a quantum critical point but in bread ranges of parameters and in different materials. It is hard to imagine that there will always be a suitable quantum critical point available in all cases. The linear temperature dependence of the resistivity can also be explained by a scattering rate that is proportional to the temperature. An appealing hypothesis, known as “Planckian dissipation” [22], states that the scattering rate is universal and given just $k_{B}T/\hbar$ where $k_{B}$ is the Boltzmann constant and $\hbar$ is the Planck constant. This picture would suggest that in the strange metal enjoys some sort of scale invariance up to arbitrarily high temperatures. While this scenario is natural in the AdS/CFT picture, it is unclear why it should apply. Recent work on generalizations of the SYK model have clarified the connection between the AdS/CFT theories and strange metals [23].
@@ IV Electronic Liquid-Crystal Phases
As we said before, high temperature superconductors arise from doping a material which is a Mott insulator. In the copper-oxide superconductors the Mott insulator is a Néel antiferromagnet with a critical temperature $T_{N}\sim 300$K with ordering wave vector $\bm{Q}=(\pi,\pi)$. In 1995 John Tranquada and coworkers discovered that the copper oxide La1.6-xNd0.4SrxCuO4 at doping near $x=1/8$ exhibits static incommensurate spin order with unidirectional order with wave vector $\bm{Q}(\delta)=(\pi-\delta,\pi-\delta)$ with an associated periodic modulation of the local charge distribution with wave vector $\bm{K}=2\bm{Q}(\delta)$ [28, 36]. In La1.6-xNd0.4SrxCuO4 the charge order has period $4$ and was consistent with charges accumulating at anti-phase domain walls of the spin order. It took more than ten years to understand the wider significance of this discovery.
@@
On the other hand, Congjun Wu, Kai Sun, Shoucheng Zhang and I showed that a Pomeranchuk instability in the spin triplet channel leads to generalizations of the nematic phase in the spin channel [47]. One example is the “nematic-spin-nematic” phase, conjectured in Ref.[33] and shown in Fig.7. This state is invariant under a $\pi/2$ spatial rotation followed by a spin flip. This state has the same symmetry of the altermagnet states which are being widely discussed in the current literature [48]. In Ref.[47] it is also shown that there is a second class of anisotropic phases in the triplet channel in which the spin polarization winds twice (for $\ell=2$) around the Fermi surface with a non-vanishing Berry phase.
@@
The order parameter of a two-dimensional nematic is any $2\times 2$ traceless tensor $\mathcal{Q}$
@@
Here $1$ and $2$ represent two orthogonal directions, say $x$ and $y$. A tensor of this type transforms as a quadrupole and is invariant under rotations by $\pi$. One can also use the two non-vanishing components of the tensor $\mathcal{Q}$ to for a director: a vector with an orientation but not a direction. There are many choices for this tensor. Phenomenologically one can use any physical observable which transforms properly under rotations such as the resistivity tensor
@@ VI.1 Intertwined Orders in La2-xBaxCuO4
La2-xBaxCuO4 has a complex structure. At relatively high temperatures the crystal structure is in the low temperature orthorhombic (LTO) phase and there is a doping-dependent structural transition, with a broad maximum for $x\sim 1/8$ at about $52$ K, to the low temperature tetragonal (LTT) phase. In the LTT phase the copper oxide planes form pairs in which each plane is orthorhombic but rotated by $90^{\circ}$ from the other plane in the pair. In the LTT phase of La2-xBaxCuO4 there is static charge stripe order (denoted by CO in Fig.9) which is unidirectional in each plane, with periodicity of $4$ lattice constants, and rotated by $90^{\circ}$ between planes [74, 29]. In other materials of the Lanthanum series of the cuprates, e.g. in La1.8-xEu0.2SrxCuO4 , the LTO-LTT structural transition happens at much higher temperatures than the charge ordering transition (see Ref.[75].) In the LTO phase the charge stripe order is “fluctuating” (in the sense of Ref.[33]) seen only in the inelastic channel at low frequencies and with a doping-dependent ordering wave vector [76]. At $x\sim 1/8$ a cascade of phase transitions are seen below $52$K where static charge stripe order is seen. Then at $\sim 42$K there is a transition a phase with static stripe unidirectional spin order with ordering wave vector $\bm{Q}_{\rm SDW}=2\bm{Q}_{\rm CDW}$. Static charge and spin orders are commensurate with each other.
@@ VI.2 The Pair-Density Wave State
How does this state explain the observed anomalies and, in particular, the dynamical layer decoupling? First the rotation of consecutive planes suppresses the inter-layer Josephson effect (pair-tunneling) since locally the superconducting orders are phase shifted by $\pi$. Notice that it does not suppress the double Josephson effects, i.e. tunneling of quartets. In addition the spins are also rotated by $\pi/2$ (dues to spin-orbit effects). Finally, Coulomb repulsion forces the unit cell to have period 4 along the c-axis. Thus, the planes with an allowed Josephson coupling are extremely far from each other and the effective coupling is essentially negligible. The result is that the Josephson effect is frustrated and the planes are effectively decoupled. Away from 1/8 doping, defects, such as local charge disorder, and dislocations of the charge order lead to a small but finite Josephson coupling between nearby planes and favor the uniform superconducting order and the suppression of the dynamical layer decoupling.
@@ VI.3 Landau Theory Picture of the PDW State
The representation of Eq.(15) describes different phases: a uniform superconductor is $\langle\Delta_{0}\rangle\neq 0$ and $\langle\Delta_{\bm{Q}}\rangle=\langle\Delta_{-\bm{Q}}\rangle=0$, b) a PDW if $\langle\Delta_{0}\rangle=0$ and $\langle\Delta_{\rm Q}\rangle=\langle\Delta_{-\rm Q}\rangle\neq 0$, c) an inversion-breaking superconductor if $\langle\Delta_{0}\rangle=0$, $\langle\Delta_{\rm Q}\rangle\neq 0$ and $\langle\Delta_{-\rm Q}\rangle=0$, and d) a striped superconductor if $\langle\Delta_{0}\rangle\neq 0$, $\langle\Delta_{\bm{Q}}\rangle\neq 0$ and $\langle\Delta_{-\bm{Q}}\rangle\neq 0$. Each one of these states has distinct symmetry properties. The inversion-breaking superconductor was predicted by Fulde and Ferrell [90] (FF) as arising in the presence of a magnetic field much in the same way as the state proposed by Larkin and Ovchinnikov. The LO and FF states are commonly referred together as the FFLO states (even though they have different symmetry) and the there is a robust literature on them as magnetic-driven superconducting states [91]. Instead, here we are interested in these translation symmetry breaking superconducting states as arising in the absence of an external magnetic field and having periodicities of a few lattice constants. On the other hand, with the exception of the Lanthanum family of the coper oxide superconductors, most states normally referred to as a PDW in the literature are actually striped superconductors. We will see below why this distinction matters.
@@
For $v_{\rm PDW}>0$ either $\Delta){\bm{Q}}\neq 0$ and $\Delta_{-\bm{Q}}=0$ (or viceversa) this is an FF state. For $v_{\rm PDW}<0$, $\Delta_{\bm{Q}}=\Delta_{-\bm{Q}}\neq 0$ and this is the PDW state. For simplicity I ignored that the stiffness of a unidirectional PDW $\kappa_{\rm PDW}$ is anisotropic .
@@
The other consequence of Eq.(21) is that if $\langle\Delta_{0}\rangle\neq 0$ and $\langle\Delta_{\bm{Q}}\rangle\neq 0$ then $\langle\rho_{\bm{K}}\rangle\neq 0$. Thus if both the uniform component and the PDW component are present then a CDW order is induced with same same ordering wave vector $\bm{Q}$ as the PDW, in addition to the $2\bm{Q}$ CDW order already present in the PDW state. Alternatively we can say that if $\langle\Delta_{0}\rangle\neq 0$ and $\langle\Delta_{\bm{Q}}\rangle\neq 0$ then there will be an induced $\bm{K}=\bm{Q}$ component (usually referred to as the“$1\bm{Q}$” component) in the CDW order with $\rho_{\bm{Q}}\sim\Delta_{0}^{*}\Delta_{\bm{Q}}$.
@@ VI.4 Topological excitations of the PDW superconductor
Figure 14: Structure of the half vortex of a PDW: at the core of the half-vortex $\lim_{r\to 0}\Delta_{\bm{Q}}=0$ but $\lim_{r\to 0}\Delta_{-\bm{Q}}\neq 0$ (and actually it increases from its value at long distances where $\lim_{r\to\infty}|\Delta_{\bm{Q}}|=\lim_{r\to\infty}|\Delta_{-\bm{Q}}|=\Delta_{\rm PDW}$). Hence, the core of the PDW vortex has an “FF halo” which breaks inversion symmetry. In this halo the $2\bm{Q}$ CDW charge order is suppressed, $\rho_{2\bm{Q}}\to 0$ as is the induced charge $4e$ superconducting component, $\Delta_{4e}\to 0$ (from Ref.[106]).
@@ VI.5 Thermal Phase Transitions and the Role of Dimensionality
In addition, there is also the problem of understanding the role of composite orders and whether there are phases characterized by composite orders, vestigial orders in the sense of Ref.[49] (for a review see Ref.[60]), exist after the “mother” PDW phase has melted. This problem was considered long ago by Golubović and Kostić [117] in a suitable large-$N$ limit of Landau-Ginzburg type theories with several order parameters. Again, this approach is highly restricted to the close proximity of a rather complex multi-critical point.
@@
where the coupling constants $g_{1,0}$, $g_{0,1}$ and $g_{1/2,1/2}$ are, respectively, proportional to the logarithm of the fugacities of the vortex operator, $\textrm{exp}(\pm i\pi\vartheta)$, (with topological charge $(1,0)$), the double dislocation operator, $\textrm{exp}(\pm 2\pi i\varphi)$, (with topological charge $(0,1)$) and the half-vortex operators, $\textrm{exp}(i\pi\vartheta\pm i\pi\varphi)$ (with topological charge (1/2/,1/2)(1/2/,1/2)( 1 / 2 / , 1 / 2 ). These three topological excitations have an energy cost which scales withe logarithm of the system size. The Kosterlitz-Thouless energy-entropy criterion then yields the critical temperatures. However these phase transitions may not necessarily happen at the same temperature.
@@
The result of this analysis is presented in Fig.16. The straight lines where these operators are marginal are generally different. In the figure the half-vortex operator (±$(\pm 1/2,\mp 1/2)$ becomes marginal at the red line, the vortex operator (±$(\pm 1,0)$ becomes marginal at the blue line, and the double dislocation operator $(0,\pm 1)$ becomes marginal at the orange line. Here the incommensurate PDW is the low temperature phase fo all values of $\kappa/\rho_{s}$. For values of $\kappa/\rho_{s}$ to the right of the point denoted by $P^{\prime}$, with increasing temperature the blue line is reached, where the vortices become relevant, the superfluidity is lost. Here the system enters an (again critical) non superfluid phase with unidirectional incommensurate CDW state (with wave vector $2\bm{Q}$). Eventually, the CDW order is melted at a higher temperature phase transition at the red line. This high temperature phase is uniform and anisotropic. This is a Nematic phase which at higher temperatures melts, at an Ising-type phase transition, into a uniform and isotropic normal phase. On the other hand, for $\kappa/\rho_{s}$ between the points $P$ and $P^{\prime}$ the PDW melts into the high temperature Nematic phase by proliferation of the half-vortices. Finally, for $\kappa/\rho_{s}$ to the left of $P$, the double dislocations proliferate at the orange line and the system enters a uniform phase which is a charge $4e$ condensate, which in turn melts into the Nematic phase by proliferating the half-vortices.
@@
The CDW phase and the charge $4e$ superconducting phase are examples of vestigial phases characterized by the composite operators defined in Eq.(18) and Eq.(19), in which some of the (almost) broken symmetries of the PDW are restored. While the CDW phase is a non-superconducting state with translation symmetry broken in only one direction (with wave vector $2\bm{Q}$), the charge $4e$ superconducting phase is exotic in the sense that its fundamental vortices are the half-vortices. All these phase transitions are very similar to the conventional Kosterlitz-Thouless (KT) transition. Except for the direct PDW-Nematic transition, at which the full $U(1)\times U(1)$ is (almost) spontaneously broken, the other transitions are conventional KT transitions at which only one of the $U(1)$ symmetries is restored. In Fig.16 , the points $P$ and $P^{\prime}$ are multi-critical points at which the $U(1)\times U(1)$ symmetry is enhanced (dynamically) and are described by an $SU(3)_{1}$ conformal field theory [123].

### v2

@@ II.1 Landau Fermi Liquids
The pole structure of the propagator implies that in the Fermi liquid there is a Fermi surface at $|\bm{p}|=k_{F}$ (here for simplicity we assumed an isotropic system) where the quasiparticles are long-lived. The analytic structure of the propagator also implies, among other things, that the resistivity has a $T^{2}$ temperature dependence  [^1] Naturally, the resistivity is finite only in the presence of momentum (and/or energy) relaxation mechanisms such as umklapp processes, disorder, phonons, etc. The temperature dependence of the resistivity is different for different processes. The $T^{2}$ dependence is for a Fermi liquid with some weak relaxation mechanism, see Ref. [4].. The existence of the Fermi surface is documented in different types of experimental probes such as quantum oscillations of resistivity, magnetization and electronic specific heat in magnetic fields, and in angle-resolved photoemission (ARPES) experiments.
@@ II.2 Ginzburg-Landau Description of the Superconducting State
The factor of 222 in Eqs.(2), (3) and (4) anticipates the fact the superconducting order parameter is a condensate of Cooper pairs and, hence, has to transform as a charge 222 complex field.
@@ II.3 BCS Theory
where $\hat{\Delta}(\bm{r})$ is the Fourier transform of the pair field operator of Eq.(5). This condition, known as off-diagonal long range order, implies that in the superconducting state the global $U(1)$ symmetry is spontaneously broken and that electric charge is conserved only modulo two.
@@ III High Temperature Superconductors are Different
[…]
@@
To this date there isn’t good theory of the strange metal phase and its explanation remains one of the challenges to Condensed Matter Theory. A commonly used theoretical proposal is to assume that there is a quantum critical point somewhere under the superconducting dome. Although this is an appealing proposal it has problems. One is that it is hard to find a quantum critical theory which naturally explains the linear resistivity. The other problem is that this behavior is seen not just in close proximity to a quantum critical point but in bread ranges of parameters and in different materials. It is hard to imagine that there will always be a suitable quantum critical point available in all cases. The linear temperature dependence of the resistivity can also be explained by a scattering rate that is proportional to the temperature. An appealing hypothesis, known as “Planckian dissipation” [23], states that the scattering rate is universal and given just $k_{B}T/\hbar$ where $k_{B}$ is the Boltzmann constant and $\hbar$ is the Planck constant. This picture would suggest that in the strange metal enjoys some sort of scale invariance up to arbitrarily high temperatures. While this scenario is natural in the AdS/CFT picture, it is unclear why it should apply. Recent work on generalizations of the SYK model have clarified the connection between the AdS/CFT theories and strange metals (for details see Ref. [24]).
@@ IV Electronic Liquid-Crystal Phases
As we said before, high temperature superconductors arise from doping a material which is a Mott insulator. In the copper-oxide superconductors the Mott insulator is a Néel antiferromagnet with a critical temperature $T_{N}\sim 300$K with ordering wave vector $\bm{Q}=(\pi,\pi)$. In 1995 John Tranquada and coworkers discovered that the copper oxide La1.6-xNd0.4SrxCuO4 at doping near $x=1/8$ exhibits static incommensurate spin order with unidirectional order with wave vector $\bm{Q}(\delta)=(\pi-\delta,\pi-\delta)$ with an associated periodic modulation of the local charge distribution with wave vector $\bm{K}=2\bm{Q}(\delta)$ [29, 37]. In La1.6-xNd0.4SrxCuO4 the charge order has period 444 and was consistent with charges accumulating at anti-phase domain walls of the spin order. It took more than ten years to understand the wider significance of this discovery.
@@
On the other hand, Congjun Wu, Kai Sun, Shoucheng Zhang and I showed that a Pomeranchuk instability in the spin triplet channel leads to generalizations of the nematic phase in the spin channel [48]. One example is the “nematic-spin-nematic” phase, conjectured in Ref.[34] and shown in Fig.7. This state has d-wave symmetry: is invariant under a $\pi/2$ spatial rotation followed by a spin flip. This state has the same symmetry of the altermagnet states which are being widely discussed in the current literature [49]. Much as its charge cousin, these spin triplet particle-hole condensates at zero total momentum cab also have higher angular momentum. In Ref.[48] it is also shown that there is a second class (the “beta-phase” in the terminology of Ref. [48]) of anisotropic phases in the triplet channel in which the spin polarization winds twice (for $\ell=2$) around the Fermi surface with a non-vanishing Berry phase. The Pomeranchuk instability is one of the possible mechanisms for the formation of these states (just as it is for the nematic) but it is natural in the continuum description of the Landau Fermi liquid. However, in most materials that display altermagnetism the lattice (and its symmetries) plays a central role and other mechanisms are possible as well. Nevertheless all altermagnets have the same symmetry breaking as the d-wave spin triplet state of Ref.[48] and, in this sense, they are the same state with the caveat that many altermagnetic materials are insulating whereas the Pomeranchuk counterpart is a metal. A discussion of the similarities and differences can be found in Ref.[50].
@@
The order parameter of a two-dimensional nematic is any $\times 2$ 2 traceless tensor $\mathcal{Q}$
@@
Here 111 and 222 represent two orthogonal directions, say $x$ and $y$. A tensor of this type transforms as a quadrupole and is invariant under rotations by $\pi$. One can also use the two non-vanishing components of the tensor $\mathcal{Q}$ to for a director: a vector with an orientation but not a direction. There are many choices for this tensor. Phenomenologically one can use any physical observable which transforms properly under rotations such as the resistivity tensor
@@ VI.1 Intertwined Orders in La2-xBaxCuO4
La2-xBaxCuO4 has a complex structure. At relatively high temperatures the crystal structure is in the low temperature orthorhombic (LTO) phase and there is a doping-dependent structural transition, with a broad maximum for $x\sim 1/8$ at about 525252 K, to the low temperature tetragonal (LTT) phase. In the LTT phase the copper oxide planes form pairs in which each plane is orthorhombic but rotated by $90^{\circ}$ from the other plane in the pair. In the LTT phase of La2-xBaxCuO4 there is static charge stripe order (denoted by CO in Fig.9) which is unidirectional in each plane, with periodicity of 444 lattice constants, and rotated by $90^{\circ}$ between planes [76, 30]. In other materials of the Lanthanum series of the cuprates, e.g. in La1.8-xEu0.2SrxCuO4 , the LTO-LTT structural transition happens at much higher temperatures than the charge ordering transition (see Ref.[77].) In the LTO phase the charge stripe order is “fluctuating” (in the sense of Ref.[34]) seen only in the inelastic channel at low frequencies and with a doping-dependent ordering wave vector [78]. At $x\sim 1/8$ a cascade of phase transitions are seen below  [^52] K where static charge stripe order is seen. Then at 42$\sim$ 42∼ 42K there is a transition a phase with static stripe unidirectional spin order with ordering wave vector $\bm{Q}_{\rm SDW}=2\bm{Q}_{\rm CDW}$. Static charge and spin orders are commensurate with each other.
@@ VI.2 The Pair-Density Wave State
How does this state explain the observed anomalies and, in particular, the dynamical layer decoupling? First the rotation of consecutive planes suppresses the inter-layer Josephson effect (pair-tunneling) since locally the superconducting orders are phase shifted by $\pi$. Notice that it does not suppress the double Josephson effect, i.e. tunneling of quartets. In addition the spins are also rotated by $\pi/2$ (dues to spin-orbit effects). Finally, Coulomb repulsion forces the unit cell to have period 4 along the c-axis. Thus, the planes with an allowed Josephson coupling are extremely far from each other and the effective coupling is essentially negligible. The result is that the Josephson effect is frustrated and the planes are effectively decoupled. Away from 1/8 doping, defects, such as local charge disorder, and dislocations of the charge order lead to a small but finite Josephson coupling between nearby planes and favor the uniform superconducting order and the suppression of the dynamical layer decoupling.
@@ VI.3 Landau Theory Picture of the PDW State
The representation of Eq.(15) describes different phases: a uniform superconductor is $\langle\Delta_{0}\rangle\neq 0$ and $\langle\Delta_{\bm{Q}}\rangle=\langle\Delta_{-\bm{Q}}\rangle=0$, b) a PDW if $\langle\Delta_{0}\rangle=0$ and $\langle\Delta_{\rm Q}\rangle=\langle\Delta_{-\rm Q}\rangle\neq 0$, c) an inversion-breaking superconductor if $\langle\Delta_{0}\rangle=0$, $\langle\Delta_{\rm Q}\rangle\neq 0$ and $\langle\Delta_{-\rm Q}\rangle=0$, and d) a striped superconductor if $\langle\Delta_{0}\rangle\neq 0$, $\langle\Delta_{\bm{Q}}\rangle\neq 0$ and $\langle\Delta_{-\bm{Q}}\rangle\neq 0$. Each one of these states has distinct symmetry properties. The inversion-breaking superconductor was predicted by Fulde and Ferrell [92] (FF) as arising in the presence of a magnetic field much in the same way as the state proposed by Larkin and Ovchinnikov. The LO and FF states are commonly referred together as the FFLO states (even though they have different symmetry) and the there is a robust literature on them as magnetic-driven superconducting states [93] or as arising from a Lifshitz transition at which the superfluid stiffness becomes negative. For these reasons the FFLO states have very long wavelengths, typically much longer than the lattice spacing of a material. Instead, here we are interested in these translation symmetry breaking superconducting states as arising in the absence of an external magnetic field and having periodicities of a few lattice constants. On the other hand, with the exception of the Lanthanum family of the coper oxide superconductors, most states normally referred to as a PDW in the literature are actually striped superconductors. We will see below why this distinction matters.
@@
For $v_{\rm PDW}>0$ either $\Delta_{\bm{Q}}\neq 0$ and $\Delta_{-\bm{Q}}=0$ (or viceversa) this is an FF state. For $v_{\rm PDW}<0$, $\Delta_{\bm{Q}}=\Delta_{-\bm{Q}}\neq 0$ and this is the PDW state. For simplicity I ignored that the stiffness of a unidirectional PDW $\kappa_{\rm PDW}$ is anisotropic .
@@
[…]
@@ VI.4 Topological excitations of the PDW superconductor
[…]
@@ VI.5 Thermal Phase Transitions and the Role of Dimensionality
[…]
@@
where the coupling constants $g_{1,0}$, $g_{0,1}$ and $g_{1/2,1/2}$ are, respectively, proportional to the logarithm of the fugacities of the vortex operator, $\textrm{exp}(\pm i\pi\vartheta)$, (with topological charge (1,0)(1,0)( 1 , 0 )), the double dislocation operator, $\textrm{exp}(\pm 2\pi i\varphi)$, (with topological charge (0,1)(0,1)( 0 , 1 )) and the half-vortex operators, $\textrm{exp}(i\pi\vartheta\pm i\pi\varphi)$ (with topological charge (1/2/,1/2)(1/2/,1/2)( 1 / 2 / , 1 / 2 ). These three topological excitations have an energy cost which scales withe logarithm of the system size. The Kosterlitz-Thouless energy-entropy criterion then yields the critical temperatures. However these phase transitions may not necessarily happen at the same temperature.
@@
[…]
@@
[…]